OpenStudy (anonymous):

Correct ? Which polynomial is prime? A. 4x2 - 17x + 15 B. 4x2 + 13x C. 8x2 + 7x - 16 D. x2 - 4 B.

Directrix (directrix):

Prime would mean it will not factor over some set, I assume the set of Rational numbers. B does factor: 4x2 + 13x = x ( 4x + 13) so look again.

OpenStudy (anonymous):

A. i think

Directrix (directrix):

That was my next choice but 4x^2 - 17x + 15 = (x - 3) (4x - 5) so it's something else.

OpenStudy (anonymous):

D.

Directrix (directrix):

x^2 - 4 = (x) ^2 - (2) ^2 which is in the pattern for factoring the difference of two squares so that's not it.

OpenStudy (anonymous):

C.

Directrix (directrix):

Yes. It has irrational roots.